# Generative Medical Event Models Improve with Scale

### v1

@@ 5.5 Inference Details
Inference with CoMET was performed by generating future medical event sequences via Monte Carlo sampling. For a given inference case, a patient history up to a specific moment was given as context to the model, and the model generated a specified number $d$ of output tokens $n$ times at a temperature of 1, where $d$ and $n$ depended on the needs of the individual task. If the amount of patient history exceeded the size of the model’s context window (8,192 tokens for all CoMET models), the history is left-truncated as necessary. When this is necessary, the model loses the ability to see demographics and older information, but this aligns with how the model sees truncated sequences during pretraining.
@@ 5.6 Evaluation Details
All evaluations were performed using patients from the held-out test set of patients.
@@ 5.6.1 Plausibility statistics
We selected 20,000 random patients and performed generations starting from the end of the patient’s last encounter that ended prior to 2022. To select for patients with some minimal degree of activity within health systems contributing data to Cosmos, we required that these patients have at least one prior encounter as well as one encounter beyond a year’s time from the generation start. We used $n=25$ generations for each patient, using up to 2000 tokens to reach one year’s time, and discarding those generations that did not reach time. We used these generations to measure the percent of multi-token events that were invalid (Appendix C), the prevalence and co-occurrence rates (Appendix D), and the probability calibrations of the number and types of encounters (Section 2.2.2). For prevalence and co-occurrence, we measured the overall agreement between ground truth and predicted rates using root-mean squared log error (RMSLE):
@@ 5.6.3 Disease-Specific Outcomes
Each sample provided to CoMET includes a patient’s health timeline from the first token up to and including the order of a medication indicated for one of the following conditions: T2DM, hypertension, and hyperlipidemia. When a sample exceeded the available context or left too little room for generation, we applied left truncation, removing tokens from the beginning of the sample. Given a sample, CoMET generated up to 2,000 new tokens from which predictions were made programmatically.
Binary predictions, such as determining whether an adverse event occurred within a determined time $t$, were made by scanning the generation up to time $t$ for a set of ICD-10-CM codes that describe an adverse outcome. Continuous predictions of lab values were made by uniformly sampling from within the range of values encapsulated by the relevant lab’s quantile bucket. If multiple labs of the desired type were produced in a given patient simulation, the average of the uniform samples was used. For all prediction types, generations that did not reach the time threshold $t$ required by outcome definitions were discarded. CoMET was evaluated on 30,000 samples per condition and predictions extracted from the model’s generations were micro-averaged over 80 generations per sample prior to calculating metrics.
@@ 5.6.4 Incident Disease Risk Prediction
All tasks here were formulated as binary prediction of a patient disease state within a 2 year window. Each prediction point in this cohort is unique to one patient and was chosen to be the last instant of a randomly selected outpatient encounter between 2020 and 2022. All patients were required to have at least 2 encounters of either outpatient or “emergent” encounter types in the 2 years prior to the prediction date. “Emergent” encounters are inpatient admission, emergency department, and urgent care visits. Additionally, all patients were required to be at least 18 and less than 120 years old at the prediction date to be included in these cohorts. For each incident disease prediction tasks, patients are excluded if they have had any of the diagnosis codes of any type from the phenotype definition before the prediction point.
Each target’s phenotype was classified as “chronic” or “emergent” and defined by a list of ICD-10-CM codes within certain encounter types (see Table 24). A chronic target was marked positive if a diagnosis appeared in at least two outpatient encounters or in a single emergent encounter in the patient’s history. An emergent target was marked positive only if a diagnosis occurred during an emergent encounter. Only encounter diagnoses and billing diagnoses were used for this target gathering. Patients in the ”Dementia (2yr)” cohort were also required to be 60 years or older at the prediction date to make the task more difficult.
@@
Inference on these tasks followed our typical strategy with the following modifications. We generated $n=60$ simulations for each patient initially with 2,000 tokens, and we retried up to 2 times to try to get the patient generations to the 2-year prediction time duration. When evaluating patient timelines generated by CoMET, the timeline would be evaluated to have the target if the timeline has any occurrence of any ICD-10-CM codes from the target phenotype within the corresponding encounter types for that phenotype.
@@ 5.6.5 Acute on Chronic Event Prediction
The non-task specific inclusion criteria and the prediction date selection criteria used for incident disease risk prediction were also used as the base criteria for acute-on-chronic event prediction, with additional task-specific inclusion criteria applied for each condition. Each acute-on-chronic event prediction task consists of a chronic and an acute (”emergent”) phenotype, such as sickle cell disease and sickle cell crisis respectively. Find the full list of phenotypes in Table 24).
@@
To meet the acute phenotype criteria, a patient must have an emergent encounter with an encounter or billing diagnosis code within the prediction window in their ground truth data. Patients who have already had this acute event prior to the prediction date are still included within the cohort. A patient timeline generated by CoMET would be evaluated to have the target if the timeline has any occurrence of any ICD-10-CM codes from the acute phenotype within an emergent encounter. The same sampling strategy, sample weighting for performance metrics, and sampling parameters as incident disease risk prediction are used in this evaluation. Table 19 has more details on dataset size and positive prevalence.
@@ 5.6.6 Differential diagnosis
For the HPB and rheumatic diagnosis clusters, we selected 9 conditions from each to represent a range of prevalence and disease types. Specific definitions of each are in Table 22. Patients were selected for evaluation based on the presence of relevant diagnosis codes in their record. We used retrospective inclusion based on the diagnosis outcomes in order to focus this evaluation on patients who have one of the target diagnoses. In order to be included in a disease cohort, the patient had to have at least two occurrences of one of the eligible ICD-10-CM codes at separate encounters (this was done to mitigate diagnoses documented as part of rule-out diagnostic tests but for which the patient did not receive additional care). Time was measured relative to the encounter containing the first occurrence of an eligible ICD-10-CM code, which was then used as the index time $t=0$. Additionally, to be included, patients were required to both have an encounter prior to one year before the index diagnosis date as well as one encounter within one year before the index diagnosis date. We then selected patients who received one of these diagnoses. Once a patient was selected, the first occurrence of each off-target diagnosis was identified for that patient. Selecting at most 9 random encounters from the year prior to $t=0$, we had CoMET generate from the end of these encounters using $n=40$ generations and up to 2,000 output tokens to predict a patient’s 18-month risk of having that diagnosis. Additionally, we also generated from the encounter header for the encounter at $t=0$ to measure CoMET’s final predictions. We chose not to generate from within this encounter due to an increased chance of information leakage and to focus the evaluation on early diagnosis detection and differential diagnosis.
@@
For the training of the task-specific supervised models, we selected many additional patients. For a proper comparison to CoMET, we chose to include patients who were being worked up for, or were at risk of developing, a much wider set of diagnoses than HPB or rheumatic diagnoses. This makes the comparison to CoMET more appropriate and also better reflects the diverse clinical scenarios a deployed system might encounter when flagging patients at risk of being diagnosed with specific diseases. For training, we expanded our pool of eligible patients by broadening the pool of possible inclusion diagnoses from those in our HPB and rheumatic lists to all 3-character ICD-10-CM codes, and selected multiple prediction dates for each as described above. We then trained 18 task-specific binary classifiers, one for each of the 18 HPB and rheumatic diagnoses.
@@
In order to visualize and contextualize the results, a patient’s predicted risk over the subsequent 18 months was used as their diagnosis risk at a moment in time, and that risk remained the same until their next encounter for which we had model predictions. In Figure 10 and Figure 11, the predictions are plotted for each day from 6 months out up to $t=0$. We first assessed CoMET’s ability to flag at-risk patients, choosing a threshold probability of 10%. Each individual plot only contains patients we had selected as having the indicated diagnosis at $t=0$. For each of the off-target lines in the plot, we also excluded patients whose ground truth record ever contained that off-target diagnosis (in the past or in the future), so that the off-target diagnosis lines represent the percent of patients flagged with an “incorrect” diagnosis.
@@ 5.6.7 Forecasting patients’ interactions with the health system
We used the same patients and CoMET generations described above in Section 5.6.1. For each encounter type indicated, we computed the mean number of encounters predicted by CoMET in the following year, as well as that predicted by the task-specific regression model.
@@ 5.6.8 Length of stay
To evaluate CoMET’s ability to predict inpatient length of stay, we randomly selected 10,000 Hospital Admissions from our test set and labeled each with the total length of stay from admission to discharge in seconds. For each hospital admission, we prompt CoMET with the patient’s history up until the point of admission (including the department specialty and any associated chief complaint and flagged primary encounter diagnosis events). We then generate up to 2,000 new tokens with $n=20$ inpatient trajectories until the encounter’s stop token appears. We take the median time-to-encounter-end as the model’s prediction.
@@
We constructed prediction targets by selecting 10,000 patients discharged between January 1, 2020, and the present date, subsequently determining whether these patients experienced a hospital readmission within 30 days. Readmissions occurring within 24 hours of discharge were excluded from this analysis because they may often reflect hospital transfers or clerically erroneous discharges [33]. A 30-day readmission in a CoMET generation was defined as the beginning of a new inpatient encounter within 30 days of the discharge time.
@@ 6 Acknowledgments
Appendix A CoMET dataset statistics
After data pre-processing as described in Section 5.1, data from Epic Cosmos was transformed into a subset used for the CoMET models, which was split into a train and test set at the patient level. Summary statistics of this dataset are shown in Table 4. Cosmos maintains and reports on over 1,000 metrics to assess data quality, focused on completeness, conformance, and plausibility. Some of these data quality metrics as measured on the CoMET dataset subset are displayed in Table 5.
@@ Appendix A CoMET dataset statistics
Table 4: CoMET dataset counts in the train, test, and full data sets, organized by various data types, demographics, and the number of source organizations contributing to each patient record. Abbreviations: BoAA = Black or African American, AI/AN = American Indian or Alaska Native, NHOPI = Native Hawaiian or Other Pacific Islander. Only breakdowns for First Race are shown.
@@
Table 5: CoMET dataset data quality metrics by data type
@@ Appendix B Tabulated results
Evaluation Task Metric CoMET-S CoMET-M CoMET-L
@@
Table 6: Percent, RMSLE, and ECE score comparisons across plausibility and encounter frequency tasks (Section 2.2.1 and Section 2.2.2) for each CoMET model.
Evaluation Task Reference CoMET-S CoMET-M CoMET-L
@@
Table 7: PR-AUC score comparisons for single-encounter generation tasks (Section 2.2.3) for each CoMET model and various baseline lookback windows.
Condition Task (time horizon) Supervised CoMET-S CoMET-M CoMET-L
@@
Table 8: AUCROC on disease-specific outcome tasks (Section 2.3.1) for each CoMET model and the best-performing supervised task-specific model.
Task (time horizon) Supervised CoMET-S CoMET-M CoMET-L
@@
Table 9: AUCROC and PR-AUC on acute-on-chronic tasks (Section 2.3.2) for each CoMET model and the best-performing supervised task-specific model.
Task (time horizon) Supervised CoMET-S CoMET-M CoMET-L
@@
Table 10: AUCROC and PR-AUC on incident disease risk prediction (Section 2.3.3) for each CoMET model and the best-performing supervised task-specific model.
Evaluation Task Supervised CoMET-S CoMET-M CoMET-L
@@
Table 11: AUCROC for differential diagnosis prediction (Section 2.4) at $t=0$ (i.e., the start of the encounter of the target diagnosis event) for each CoMET model and the best-performing supervised task-specific model.
Task (time horizon) Supervised CoMET-S CoMET-M CoMET-L
@@
Table 12: MAE for one-year encounter count forecasting (Section 2.5) for each CoMET model and the best-performing supervised task-specific model.
Task Supervised CoMET-S CoMET-M CoMET-L
@@
Table 13: MAE on length of stay tasks (Section 2.5.1) measured in days for each CoMET model and the best-performing supervised task-specific model.
@@
CoMET-L 0.770
CoMET-M 0.723
CoMET-S 0.706
@@
Table 14: AUCROC scores on the 30-day readmission task (Section 2.5.1) for the CoMET models and the best-performing task-specific model.
@@ Appendix C Multi-token event validity
Each of the CoMET models was evaluated for the frequency of generating invalid multi-token events. A diagnostic event is considered invalid if there are ICD-10-CM tokens in combination that do not correspond to real ICD-10-CM codes. The same was done for medication orders with ATC codes. A lab result is considered invalid if the lab token is not followed by a lab result quantile token. Encounter headers were considered invalid if an encounter start token was not directly followed by a department specialty token (as noted in Table 5, not all encounters have a specified department specialty, but the encounter headers still have a token with a specialty type of ’unspecified’). These experiments were done with 20,000 patients, each with 25 1-year generations. The denominator for each metric is the number of tokens that initiated an event.
Figure 19: Generation syntactic validity. The percent of diagnosis, medication, lab result, and encounter header events that were invalidly generated by each of the CoMET models.

### v3

@@ 5.5 Inference Details
Inference with Curiosity was performed by generating future medical event sequences via Monte Carlo sampling. For a given inference case, a patient history up to a specific moment was given as context to the model, and the model generated a specified number $d$ of output tokens $n$ times at a temperature of 1, where $d$ and $n$ depended on the needs of the individual task. If the amount of patient history exceeded the size of the model’s context window (8,192 tokens for all Curiosity models), the history is left-truncated as necessary. When this is necessary, the model loses the ability to see demographics and older information, but this aligns with how the model sees truncated sequences during pretraining.
@@ 5.6 Evaluation Details
All evaluations were performed using patients from the held-out test set.
@@ 5.6.1 Plausibility statistics
We selected 20,000 random patients and performed generations starting from the end of the patient’s last encounter that ended prior to 2022. To select for patients with some minimal degree of activity within health systems contributing data to Cosmos, we required that these patients have at least one prior encounter as well as one encounter beyond a year’s time from the generation start. We used $n=25$ generations for each patient, using up to 2,000 tokens to reach one year’s time, and discarding those generations that did not reach time. We used these generations to measure the percent of multi-token events that were invalid (Appendix C), the prevalence and co-occurrence rates (Appendix D), and the probability calibrations of the number and types of encounters (Section 2.2.2). For prevalence and co-occurrence, we measured the overall agreement between ground truth and predicted rates using root-mean squared log error (RMSLE):
@@ 5.6.3 Disease-Specific Outcomes
Each sample provided to Curiosity includes a patient’s health timeline from the first token up to and including the order of a medication indicated for one of the following conditions: T2DM, hypertension, and hyperlipidemia. When a sample exceeded the available context or left too little room for generation, we applied left truncation, removing tokens from the beginning of the sample. Given a sample, Curiosity generated up to 2,000 new tokens from which predictions were made programmatically.
Binary predictions, such as determining whether an adverse event occurred within a determined time $t$, were made by scanning the generation up to time $t$ for a set of ICD-10-CM codes that describe an adverse outcome. Continuous predictions of lab values were made by uniformly sampling from within the range of values encapsulated by the relevant lab’s quantile bucket. If multiple labs of the desired type were produced in a given patient simulation, the average of the uniform samples was used. For all prediction types, generations that did not reach the time threshold $t$ required by outcome definitions were discarded. Curiosity was evaluated on 30,000 samples per condition and predictions extracted from the model’s generations were micro-averaged over 80 generations per sample prior to calculating metrics.
@@ 5.6.4 Incident Disease Risk Prediction
All tasks here were formulated as binary prediction of a patient disease state within a 2-year window. Each prediction point in this cohort is unique to one patient and was chosen to be the last instant of a randomly selected outpatient encounter between 2020 and 2022. All patients were required to have at least 2 encounters of either outpatient or “emergent” encounter types in the 2 years prior to the prediction date. “Emergent” encounters are inpatient admission, emergency department, and urgent care visits. Additionally, all patients were required to be at least 18 and less than 120 years old at the prediction date to be included in these cohorts. For each incident disease prediction tasks, patients are excluded if they have had any of the diagnosis codes of any type from the phenotype definition before the prediction point.
Each target’s phenotype was classified as “chronic” or “emergent” and defined by a list of ICD-10-CM codes within certain encounter types (see Table 24). A chronic target was marked positive if a diagnosis appeared in at least two outpatient encounters or in a single emergent encounter in the patient’s history. An emergent target was marked positive only if a diagnosis occurred during an emergent encounter. Only encounter diagnoses and billing diagnoses were used for this target gathering. Patients in the “Dementia (2yr)” cohort were also required to be 60 years or older at the prediction date to make the task more difficult.
@@
Inference on these tasks followed our typical strategy with the following modifications. We generated $n=60$ simulations for each patient initially with 2,000 tokens, and we retried up to two times to get the patient generations to the 2-year prediction time duration. When evaluating patient timelines generated by Curiosity, the timeline would be evaluated to have the target if the timeline has any occurrence of any ICD-10-CM codes from the target phenotype within the corresponding encounter types for that phenotype.
@@ 5.6.5 Acute on Chronic Event Prediction
The non-task specific inclusion criteria and the prediction date selection criteria used for incident disease risk prediction were also used as the base criteria for acute-on-chronic event prediction, with additional task-specific inclusion criteria applied for each condition. Each acute-on-chronic event prediction task consists of a chronic and an acute (“emergent”) phenotype, such as sickle cell disease and sickle cell crisis respectively. Find the full list of phenotypes in Table 24).
@@
To meet the acute phenotype criteria, a patient must have an emergent encounter with an encounter or billing diagnosis code within the prediction window in their ground truth data. Patients who have already had this acute event prior to the prediction date are still included within the cohort. A patient timeline generated by Curiosity would be evaluated to have the target if the timeline has any occurrence of any ICD-10-CM codes from the acute phenotype within an emergent encounter. The same sampling strategy, sample weighting for performance metrics, and sampling parameters as incident disease risk prediction are used in this evaluation. Table 19 has more details on dataset size and positive prevalence.
@@ 5.6.6 Differential diagnosis
For the HPB and rheumatic diagnosis clusters, we selected 9 conditions from each to represent a range of prevalence and disease types. Specific definitions of each are in Table 22. Patients were selected for evaluation based on the presence of relevant diagnosis codes in their record. We used retrospective inclusion based on the diagnosis outcomes in order to focus this evaluation on patients who have one of the target diagnoses. In order to be included in a disease cohort, the patient had to have at least two occurrences of one of the eligible ICD-10-CM codes at separate encounters (this was done to mitigate diagnoses documented as part of rule-out diagnostic tests but for which the patient did not receive additional care). Time was measured relative to the encounter containing the first occurrence of an eligible ICD-10-CM code, which was then used as the index time $t=0$. Additionally, to be included, patients were required to both have an encounter prior to one year before the index diagnosis date as well as one encounter within one year before the index diagnosis date. We then selected patients who received one of these diagnoses. Once a patient was selected, the first occurrence of each off-target diagnosis was identified for that patient. Selecting at most 9 random encounters from the year prior to $t=0$, we had Curiosity generate from the end of these encounters using $n=40$ generations and up to 2,000 output tokens to predict a patient’s 18-month risk of having that diagnosis. Additionally, we also generated from the encounter header for the encounter at $t=0$ to measure Curiosity’s final predictions. We chose not to generate from within this encounter due to an increased chance of information leakage and to focus the evaluation on early diagnosis detection and differential diagnosis.
@@
For the training of the task-specific supervised models, we selected many additional patients. For a proper comparison to Curiosity, we chose to include patients who were being worked up for, or were at risk of developing, a much wider set of diagnoses than HPB or rheumatic diagnoses. This makes the comparison to Curiosity more appropriate and also better reflects the diverse clinical scenarios a deployed system might encounter when flagging patients at risk of being diagnosed with specific diseases. For training, we expanded our pool of eligible patients by broadening the pool of possible inclusion diagnoses from those in our HPB and rheumatic lists to all 3-character ICD-10-CM codes, and selected multiple prediction dates for each as described above. We then trained 18 task-specific binary classifiers, one for each of the 18 HPB and rheumatic diagnoses.
@@
In order to visualize and contextualize the results, a patient’s predicted risk over the subsequent 18 months was used as their diagnosis risk at a moment in time, and that risk remained the same until their next encounter for which we had model predictions. In Figure 10 and Figure 11, the predictions are plotted for each day from 6 months out up to $t=0$. We first assessed Curiosity’s ability to flag at-risk patients, choosing a threshold probability of 10%. Each individual plot only contains patients we had selected as having the indicated diagnosis at $t=0$. For each of the off-target lines in the plot, we also excluded patients whose ground truth record ever contained that off-target diagnosis (in the past or in the future), so that the off-target diagnosis lines represent the percent of patients flagged with an “incorrect” diagnosis.
@@ 5.6.7 Forecasting patients’ interactions with the health system
We used the same patients and Curiosity generations described above in Section 5.6.1. For each encounter type indicated, we computed the mean number of encounters predicted by Curiosity in the following year, as well as that predicted by the task-specific regression model.
@@ 5.6.8 Length of stay
To evaluate Curiosity’s ability to predict inpatient length of stay, we randomly selected 10,000 Hospital Admissions from our test set and labeled each with the total length of stay from admission to discharge in seconds. For each hospital admission, we prompted Curiosity with the patient’s history up until the point of admission (including the department specialty and any associated chief complaint and flagged primary encounter diagnosis events). We then generated up to 2,000 new tokens with $n=20$ inpatient trajectories until the encounter’s stop token appears. We took the median time-to-encounter-end as the model’s prediction.
@@
We constructed prediction targets by selecting 10,000 patients discharged between January 2020 and March 2025, subsequently determining whether these patients experienced a hospital readmission within 30 days. Readmissions occurring within 24 hours of discharge were excluded from this analysis because they may often reflect hospital transfers or clerically erroneous discharges [33]. A 30-day readmission in a Curiosity generation was defined as the beginning of a new inpatient encounter within 30 days of the discharge time.
@@ 6 Acknowledgments
Appendix A Curiosity dataset statistics
After data pre-processing as described in Section 5.1, data from Epic Cosmos was transformed into a subset used for the Curiosity models, which was split into a train and test set at the patient level. Summary statistics of this dataset are shown in Table 4. Cosmos maintains and reports on over 1,000 metrics to assess data quality, focused on completeness, conformance, and plausibility. Some of these data quality metrics as measured on the Curiosity dataset subset are displayed in Table 5.
@@ Appendix A Curiosity dataset statistics
Table 4: Curiosity dataset counts in the train, test, and full data sets, organized by various data types, demographics, and the number of source organizations contributing to each patient record. Abbreviations: BoAA = Black or African American, AI/AN = American Indian or Alaska Native, NHOPI = Native Hawaiian or Other Pacific Islander. Only breakdowns for First Race are shown.
@@
Table 5: Curiosity dataset data quality metrics by data type
@@ Appendix B Tabulated results
Evaluation Task Metric Curiosity-S Curiosity-M Curiosity-L
@@
Table 6: Percent, RMSLE, and ECE score comparisons across plausibility and encounter frequency tasks (Section 2.2.1 and Section 2.2.2) for each Curiosity model.
Evaluation Task Reference Curiosity-S Curiosity-M Curiosity-L
@@
Table 7: PR-AUC score comparisons for single-encounter generation tasks (Section 2.2.3) for each Curiosity model and various baseline lookback windows.
Condition Task (time horizon) Supervised Curiosity-S Curiosity-M Curiosity-L
@@
Table 8: AUCROC on disease-specific outcome tasks (Section 2.3.1) for each Curiosity model and the best-performing supervised task-specific model.
Task (time horizon) Supervised Curiosity-S Curiosity-M Curiosity-L
@@
Table 9: AUCROC and PR-AUC on acute-on-chronic tasks (Section 2.3.2) for each Curiosity model and the best-performing supervised task-specific model.
Task (time horizon) Supervised Curiosity-S Curiosity-M Curiosity-L
@@
Table 10: AUCROC and PR-AUC on incident disease risk prediction (Section 2.3.3) for each Curiosity model and the best-performing supervised task-specific model.
Evaluation Task Supervised Curiosity-S Curiosity-M Curiosity-L
@@
Table 11: AUCROC for differential diagnosis prediction (Section 2.4) at $t=0$ (i.e., the start of the encounter of the target diagnosis event) for each Curiosity model and the best-performing supervised task-specific model.
Task (time horizon) Supervised Curiosity-S Curiosity-M Curiosity-L
@@
Table 12: MAE for one-year encounter count forecasting (Section 2.5) for each Curiosity model and the best-performing supervised task-specific model.
Task Supervised Curiosity-S Curiosity-M Curiosity-L
@@
Table 13: MAE on length of stay tasks (Section 2.5.1) measured in days for each Curiosity model and the best-performing supervised task-specific model.
@@
Curiosity-L 0.770
Curiosity-M 0.723
Curiosity-S 0.706
@@
Table 14: AUCROC scores on the 30-day readmission task (Section 2.5.1) for the Curiosity models and the best-performing task-specific model.
@@ Appendix C Multi-token event validity
Each of the Curiosity models was evaluated for the frequency of generating invalid multi-token events. A diagnostic event is considered invalid if there are ICD-10-CM tokens in combination that do not correspond to real ICD-10-CM codes. The same was done for medication orders with ATC codes. A lab result is considered invalid if the lab token is not followed by a lab result quantile token. Encounter headers were considered invalid if an encounter start token was not directly followed by a department specialty token (as noted in Table 5, not all encounters have a specified department specialty, but the encounter headers still have a token with a specialty type of ’unspecified’). These experiments were done with 20,000 patients, each with 25 1-year generations. The denominator for each metric is the number of tokens that initiated an event.
Figure 19: Generation syntactic validity. The percent of diagnosis, medication, lab result, and encounter header events that were invalidly generated by each of the Curiosity models.